OpenStudy (anonymous):

Write the equation of the line given the slope and the y-intercept. Type your answer in the box without skipping any spaces. m = 5, y-intercept (0, -3)

OpenStudy (saifoo.khan):

y=5x-3

OpenStudy (saifoo.khan):

y = mx + b where m is slope. b is y-intercpt.

OpenStudy (anonymous):

so wat if it only gives me tha slope and pont on tha line like this Write the equation of the line given the slope and a point on the line. Type your answer in the box without skipping any spaces. m = 7, point (-2, 0)

OpenStudy (anonymous):

wat do i do then

OpenStudy (saifoo.khan):

this equation is a bit different as u have the x-intercpt. simply insert the values in y-mx+b to get the value for b.

OpenStudy (anonymous):

y=7x+0 ?

OpenStudy (saifoo.khan):

y = mx + b 0 = 7(-2) + b 0 = -14 + b b = 14. now you have the y-intercpt. And u had the slope before hand. so: y = 7x + 14

OpenStudy (anonymous):

oh ok so i just put in the x value and solve for b

OpenStudy (saifoo.khan):

Yesss.

OpenStudy (anonymous):

yaaaaaaaaaaaaay!

OpenStudy (saifoo.khan):

:DDDDDDD